(代数体の) Class number relation に関する文献 (2012 年 2 月更新版)

[1] G. L. Dirichlet : Untersuchungen über die Theorie der complexen Zahlen, Bericht Verhand. Königl. Preuß. Akad. Wiss. Jahrg. 1841, 190–194; [jpg], Akademiebibliothek; reprinted in G. Lejeune Dirichlet’s Werke. Herausgegeben auf Veranlassung der Königlich Preussischen Akademie der Wissenschaften von L. Kronecker, Erster Band mit G. Lejeune Dirichlet’s Bildniss, G. Keimer, Berlin, 1889, XXX, S. 503–508; [PDF, JPEG, txt], Gallica; [PDF], UMDL; reprinted by Chelsea Publishing Co., Bronx, New York, 1969; J. Reine Angew. Math. XXII (1841), Hft. 4, 375– 378; doi: 10.1515/crll.1841.22.375 [PDF (681 KB)];

PPN243919689_0022_log18.pdf (515,09 KB), GDZ;

[PDF], DigiZeitschriften.

[2] G. L. Dirichlet : Untersuchungen über die Theorie der complexen Zahlen, Abh. Königl. Preuss. Akad. Wiss. Physik.-math. Kl. 1841, 141–161; [jpg], Akademiebibliothek; reprinted in G. Lejeune Dirichlet’s Werke. Herausgegeben auf Veranlassung der Königlich Preussischen Akademie der Wissenschaften von L. Kronecker, Erster Band mit G. Lejeune Dirichlet’s Bildniss, G. Keimer, Berlin, 1889, XXXI, S. 509–532; [PDF, JPEG, txt], Gallica; [PDF], UMDL; reprinted by Chelsea Publishing Co., Bronx, New York, 1969.

[3] G. L. Dirichlet : Recherches sur les formes quadratiques à coefficients et à indéterminées complexes, J. Reine Angew. Math. XXIV (1842), Heft 4, 291–371; doi: 10.1515/crll.1842.24.291 [PDF (9310

KB)]; PPN243919689_0024_log32.pdf (7,94 MB), GDZ;

[PDF], DigiZeitschriften; reprinted in G. Lejeune Dirichlet’s Werke. Herausgegeben auf Veranlassung der Königlich Preussischen Akademie der Wissenschaften von L. Kronecker, Erster Band mit G. Lejeune Dirichlet’s Bildniss, G. Keimer, Berlin, 1889, XXXII, S. 533–618; [PDF, JPEG, txt], Gallica; [PDF], UMDL; reprinted by Chelsea Publishing Co., Bronx, New York, 1969.

[4] R. I. Bachmann : Zur Theorie der complexen Zahlen, J. Reine Angew. Math. LXVII (1867), Heft 2, 200– 204; doi: 10.1515/crll.1867.67.200 [PDF (682 KB)];

PPN243919689_0067_log16.pdf (491,37 KB), GDZ;

[PDF], DigiZeitschriften; ERAM 067.1746cj.

[5] D. Hilbert : Ueber den Dirichlet’schen biquadratischen Zahlkörper, Math. Ann. XLV (1894), Heft 3, 309–340; DOI: 10.1007/BF01446682 [xx1vv3j175676513.pdf

(1.6 MB)]; PPN235181684_0045_log29.pdf (1,99 MB),

GDZ; [PDF], DigiZeitschriften; reprinted in David Hilbert Gesammelte Abhandlungen, Erster Band: Zahlentheorie, J. Springer, Berlin, 1932, 5, S. 24–52;

PPN237821559_log10.pdf (3,21 MB), GDZ; reprint by

Chelsea Publshing Company, Bronx, New York, 1965; Zweite Auflage: Springer-Verlag, Berlin and New York, 1970; JFM 25.0303.01 (reviewed by Minkowski) [html, DVI, ps, PDF, TeX].

[6] G. Herglotz : Über einen Dirichlet’schen Satz, Math. Z. XII (1922), 255–261; DOI: 10.1007/BF01482079 [q6206705m568475k.pdf (294.9

KB)]; PPN266833020_0012_log20.pdf (538,26 KB),

GDZ; [PDF], DigiZeitschriften; JFM 48.0170.01 (reviewed by Hecke, E.) [html, DVI, ps, PDF, TeX].

[7] H. Nehrkorn : Über absolute Idealklassengruppen und Einheiten in algebraischen Zahlkörpern, Abh. Math. Sem. Univ. Hamburg 9 (1933), 318–334; DOI: 10.1007/BF02940658 [dk12g532034044t2.pdf (811.6 KB)]; JFM 59.0165.02 (reviewed by Neumann, B. H.) [html, DVI, ps, PDF, TeX].

[8] A. Scholz : Idealklassen und Einheiten in kubischen Körpern, Monatsh. Math. Phys. 40 (1933), no. 1, 211–222; DOI: 10.1007/BF01708865 [t42k56636737t227.pdf (641.8 KB)]; MR1550202; JFM 59.0194.03 (reviewed by Neisß, F.) [html, DVI, ps, PDF, TeX].

[9] S. Kuroda : Über den Dirichletschen Körper, J. Fac. Imp. Univ. Tokyo Sect. I [東京帝國大學理學部紀要. 第 1 類] 4 (1943), 383–406; MR0021031 (9, 12f) (reviewed by M. Krasner; reprinted in Reviews in Number Theory, volume 5A, R14-200) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0061.05901 (reviewed by Z. Suetuna) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[10] S. Kuroda : Über die Klassenzahlen algebraischer Zahlkörper, Nagoya Math. J. 1 (1950), 1–10; euclid.nmj/1118764698 [PDF (847 KB)]; MR0039759 (12, 593a) (reviewed by R. Brauer; reprinted in Reviews in Number Theory, volume 5, R26-6) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0037.16101 (reviewed by Hasse) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[11] R. Brauer : Beziehungen zwischen Klassenzahlen von Teilkörpern eines galoisschen Körpers, Math. Nachr. 4 (1950), Issue 1-6, 158–174; reprinted in Collected papers, vol. III. Finite groups, Lie groups, number theory, polynomials and equations; geometry and biography (P. Fonf and W. J. Wong, eds.), Mathematicians of Our Time, 19, MIT Press, Cambridge, Mass.-London, 1980, [59], pp. 497–513; MR0039760 (12, 593b) (reviewed by T. Nakayama; reprinted in Reviews in Number Theory, volume 5, R26-7) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0042.03801 (reviewed by H. Bergström) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[12] H. Hasse : Über die Klassenzahl abelscher Zahlkörper, Mathematische Lehrbucher und Monographien, Band I, Akademie-Verlag, Berlin, 1952, xii+190 S.; MR0049239 (14, 141a) (reprinted in Reviews in Number Theory, volume 5, G05-34) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0046.26003 (reviewed by S. Kuroda) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften; Nachdruck der ersten Aufgabe. Mit einer Einleitung von J. Martinet : Springer-Verlag, New York and Berlin, 1985, xxii+xii+190 S.; MR0842666 (87j:11122a) (reviewed by Editors) [html, PDF, DVI, PostScript, BibTeX]; Mathematische Lehrbucher und Monographien, II. Abteilung: Mathematische Monographien, Band 1, Akademie-Verlag, Berlin, 1985; MR0834791 (87j:11122b) (reviewed by Editors) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0668.12004 [html, PDF, XML, ASCII, DVI, PS, Bib-TeX], 0668.12003 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[13] T. Kubota : Über die Beziehung der Klassenzahlen der Unterkörper des bizyklischen biquadratischen Zahlkörpers, Nagoya Math. J. 6 (1953), 119–127; euclid.nmj/1118799485 [PDF (757 KB)]; MR0059960 (15, 605e) (reviewed by R. Brauer; reprinted in Reviews in Number Theory, volume 5, R16-12) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0053.21902 (reviewed by S. Kuroda) [scanned image]; [PDF], GDZ; [PDF], DigiZeitschriften.

[14] T. Kubota : Über den bizyklischen biquadratischer Zahlkörper, Nagoya Math. J. 10 (1956), 65–85; euclid.nmj/1118799770 [PDF (1406 KB)]; MR0083009 (18, 643e) (reviewed by G. Whaples; reprinted in Reviews in Number Theory, volume 5, R16-21) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0074.03001 (reviewed by H. W. Leopoldt) [scanned image].

[15] T. Honda : On absolute class fields of certain algebraic number fields, J. Reine Angew. Math. 203 (1960), Heft 1/2, 80–89; doi: 10.1515/crll.1960.203.80 [PDF

(1350 KB)]; PPN243919689_0203_log11.pdf (1,10 MB),

GDZ; [PDF], DigiZeitschriften; MR0115984 (22 #6781) (reviewed by G. Whaples; reprinted in Reviews in Number Theory, volume 5, R36-36) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0100.26802 (reviewed by K. Alber) [scanned image].

[16] T. Honda : On the absolute ideal class groups of metacyclic number fields of a certain type, Nagoya Math.

J. 17 (1960), 171–179; euclid.nmj/1118800461 [PDF (700 KB)]; MR0166189 (29 #3466) (reviewed by S. P. Demuškin; reprinted in Reviews in Number Theory, volume 5, R20-53) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0121.28003 (reviewed by M. Cugiani) [scanned image].

[17] H. Wada : On the class number and the unit group of certain algebraic number fields, J. Fac. Sci. Univ. Tokyo Sect. IA Math. [東京大学理学部紀要第 I 類 A 数学] 13 (1966), No. 2, 201–209; jfs130209.pdf (308 Kb), URI: http://hdl.handle.net/2261/6067; MR0214565 (35 #5414) (reviewed by T. Kubota; reprinted in Reviews in Number Theory, volume 5, R26-35) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0158.30103 (reviewed by A. Yokoyama) [scanned image].

[18] P. Barrucand and H. Cohn : A rational genus, class number divisibility and unit theory for pure cubic fields, J. Number Theory 2 (1970), no. 1, 7– 21; doi:10.1016/0022-314X(70)90003-X [PDF (730 K)]; MR0249398 (40 #2643) (reviewed by T. Kubota; reprinted in Reviews in Number Theory, volume 5, R16-54) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0192.40001 (Autorreferat) [scanned image].

[19] P. Barrucand and H. Cohn : Remarks on principal factors in a relative cubic field, J. Number Theory 3 (1971), no. 2, 226–231; doi:10.1016/0022-314X(71)90040-0 [PDF (600 K)]; MR0276197 (43 #1945) (reviewed by T. Kubota; reprinted in Reviews in Number Theory, volume 5, R16-48) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0218.12002 (Autorreferat) [scanned image].

[20] F. Halter-Koch : Ein Satz über Geschlechter relativzyklischer Zahlkörper von Primzahlgrad und seine Anwendung auf biquadratisch-bizyklische Körper, J. Number Theory 4 (1972), no. 2, 144–156; doi:10.1016/0022-314X(72)90060-1 [PDF (560 K)]; MR0302612 (46 #1756) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5A, R36-202) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0241.12004 (Autorreferat) [scanned image].

[21] M. Ishida : Fundamental units of certain algebraic number fields, Abh. Math. Sem. Univ. Hamburg 39 (1973), 245–250; DOI: 10.1007/BF02992833 [e860270808380h06.pdf (234.3 KB)]; MR0335469 (49 #250) (reviewed by P. Hartung; reprinted in Reviews in Number Theory, volume 5A, R16-215) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0311.12005 (reviewed by H. Yokoi) [scanned image].

[22] H. Lang : Über die Klassenzahl eines imaginären bizyklischen biquadratischen Zahlkörpers und seines reell-quadratischern Teilkörpers. I, J. Reine Angew. Math. 262/263 (1973), 18–40; doi: 10.1515/crll.1973.262-263.18 [PDF (2676 KB)];

PPN243919689_0262_0263_log6.pdf (2,30 MB),

GDZ; [PDF], DigiZeitschriften; MR0325573 (48 #3920) (reviewed by K. Lakkis; reprinted in Reviews in Number Theory, volume 5A, R26-211) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0285.12013 (reviewed by A. Yokoyama) [scanned image].

[23] H. Lang : Über die Klassenzahl eines imaginären bizyklischen biquadratischen Zahlkörpers und seines reellquadratischern Teilkörpers. II, J. Reine Angew. Math. 267 (1974), 175–178; doi: 10.1515/crll.1974.267.175

[PDF (615 KB)]; PPN243919689_0267_log16.pdf

(384,50 KB), GDZ; [PDF], DigiZeitschriften; MR0342492 (49 #7238) (reviewed by K. Lakkis; reprinted in Reviews in Number Theory, volume 5A, R26-211E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0406.12004 (reviewed by K. Lakkis) [scanned image].

[24] H. Lang und R. Schertz : Bemerkungen über die Klassenzahl eines imaginären bizyklischen biquadratischen Zahlkörpers, Abh. Math. Sem. Univ. Hamburg 42 (1974), 212–216; DOI: 10.1007/BF02993548 [v7386407578n6757.pdf (210.4 KB)]; MR0412149 (54 #276) (reviewed by T. Metsänkylä; reprinted in Reviews in Number Theory, volume 5A, R16-252) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0357.12007 (reviewed by T. Lepistö) [scanned image].

[25] T. Callahan : The 3-groups of non-Galois cubic fields I, Mathematika 21 (1974), no. 1, 72–89; doi:10.1112/S0025579300005805 [PDF]; MR0366876 (51 #3122) (reviewed by T. Lepistö; reprinted in Reviews in Number Theory, volume 5A, R36-209) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0312.12008 (summary) [scanned image].

[26] T. Callahan : The 3-groups of non-Galois cubic fields II, Mathematika 21 (1974), no. 2, 168–188; doi:10.1112/S002557930000855X [PDF]; MR0366876 (51 #3122) (reviewed by T. Lepistö; reprinted in Reviews in Number Theory, volume 5A, R36-209) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0312.12009 (from the introduction) [scanned image].

[27] C. J. Parry : Pure quartic number fields whose class numbers are even, J. Reine Angew. Math. 272 (1975), 102–112; doi: 10.1515/crll.1975.272.102 [PDF

(1132 KB)]; PPN243919689_0272_log13.pdf (867,68

KB), GDZ; [PDF], DigiZeitschriften; MR0364181 (51 #436) (reviewed by H. Cohn; reprinted in Reviews in Number Theory, volume 5A, R16-226) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0312.12006 (reviewed by S. Takahashi) [scanned image].

[28] C. J. Parry : Class number relations in pure sextic fields, J. Reine Angew. Math. 274/275 (1975), 360–375; doi: 10.1515/crll.1975.274-275.360 [PDF (1946 KB)];

PPN243919689_0274_0275_log36.pdf (1,64 MB), GDZ;

[PDF], DigiZeitschriften; MR0376614 (51 #12789) (reviewed by H. Cohn; reprinted in Reviews in Number Theory, volume 5A, R16-233) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0315.12006 (reviewed by A. Yokoyama) [scanned image].

[29] C. J. Parry : Class number relations in pure quintic fields, Symposia Mathematica, vol. XV (Convegno di Strutture in Corpi Algebrici, INDAM, Rome, 1973), Academic Press, London, 1975, PP. 475–485; MR0387241 (52 #8084) (reviewed by T. Lepistö; reprinted in Reviews in Number Theory, volume 5A, R26-229) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0331.12003 (reviewed by H. London) [scanned image].

[30] N. Moser : Unités et nombre de classes d’une extension diédrale de Q, Jornées Arithmétiques de Bordeaux (Conf., Univ. Bordeaux, Bordeaux, 1974), Astérisque Nos. 24–25, Soc. Math. France, Paris, 1975, pp. 29– 35; MR0376610 (52 #12785) (reviewed by H. Cohn; reprinted in Reviews in Number Theory, volume 5A, R22-229) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0313.12002 (reviewed by E. Brown) [scanned image].

[31] T. Callahan : Dihedral field extension of order 2p whose class numbers are multiples of p, Canad. J. Math. 28 (1976), no. 2, 429–439; DOI: 10.4153/CJM-1976-043-

x [cjm1976v28.0429-0439.pdf [853 KB]]; MR0412148

(54 #275) (reviewed by T. Lepistö; reprinted in Reviews in Number Theory, volume 5A, R26-233) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0356.12007 (introduction) [scanned image].

[32] C. J. Parry and C. D. Walter : The class numbers of pure fields of prime degree, Mathematika 23 (1976), no. 2, 220–226; doi:10.1112/S0025579300008822 [PDF]; MR0435032 (55 #7994) (reviewed by L. Washington; reprinted in Reviews in Number Theory, volume 5A, R26-241) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0364.12005 (reviewed by T. Metsänkylä) [scanned image]; correction : ibid. 24 (1977), no. 1, 133; MR0466067 (57 #5950) (reprinted in Reviews in Number Theory, volume 5A, R26-241E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0364.12006 (reviewed by T. Metsänkylä) [scanned image].

[33] C. D. Walter : A class number relation in Frobenius extensions of number fields, Mathematika 24 (1977), no. 2, 216–225; doi:10.1112/S0025579300009128 [PDF]; MR0472768 (57 #12458) (reviewed by Y. Furuta; reprinted in Reviews in Number Theory, volume 5A, R26-251) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0358.12004 (Autorreferat) [scanned image].

[34] F. Halter-Koch : Einheiten und Divisorenklassen in Galois’schen algebraischen Zahlkörpern mit Diedergruppe der Ordnung 2l für eine ungerade Primzahl l, Acta Arith. XXXIII (1977), fasc. 4, 355–364; aa3343.pdf, icm; MR0453695 (56 #11955) (reviewed by C. J. Parry; reprinted in Reviews in Number Theory, volume 5A, R26-248) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0416.12003 (reviewed by G. Gras) [scanned image].

[35] W. Jehne : Über die Einheiten- und Divisorenklassengruppe von reellen Frobeniuskörpern von Maximaltyp, Math. Z. 152 (1977), Heft 3, 223–252; DOI: 10.1007/BF01488966 [xj61074q68270382.pdf (1.3

MB)]; PPN266833020_0152_log22.pdf (2,14 MB),

GDZ; [PDF], DigiZeitschriften; MR0429828 (55 #2838) (reviewed by H. Koch; reprinted in Reviews in Number Theory, volume 5A, R26-239) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0334.12024 (Autorreferat) [scanned image], 0338.12008 [scanned image].

[36] C. Castela : Nombre de classes d’idéaux d’une extension diédrale d’un corps de nomres, C. R. Acad. Sci. Paris Sér. A-B 287 (1978), n^{∘} 7, A483–A486; MR0512085 (80c:12012) (summary; reprinted in Reviews in Number Theory, volume 5A, R26-256) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0395.12011 (reviewed by C. Parry) [scanned image].

[37] C. Castela : Nombre de classes d’idéaux d’une extension diédrale de degré 8 de Q, Séminaire de Théorie des Nombres 1977–1978, Exp. No. 5, 8 pp., CNRS, Talence,

1978; PPN320141322_0007_log10.pdf (676,21 KB),

GDZ; [PDF], DigiZeitschriften; MR0550266 (81e:12006) (reviewed by C. D. Walter; reprinted in Reviews in Number Theory, volume 5A, R26-251) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0405.12010 (reviewed by O. Femenko) [scanned image].

[38] F. Halter-Koch et N. Moser : Sur le nombre de classes de certaines extension métacycliques sur Q ou sur un corps quadratique imaginaire, J. Math. Soc. Japan [日本数学会欧文誌] 30 (1978), No. 2, 237–248; doi:10.2969/jmsj/03020237,

euclid.jmsj/1240432554  [PDF  (855  KB)];

JOI:JST.Journalarchive/jmath1948/30.237 [30_237. pdf (718 K)], Journal@rchive, JST; MR0485781 (58 #5587) (reviewed by J. Ritter; reprinted in Reviews in Number Theory, volume 5A, R20-228) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0368.12003 (Autorreferat) [scanned image].

[39] C. J. Parry : On the class number of relative quadratic fields, Math. Comp. 32 (1978), no. 144, 1261–1270; DOI: 10.1090/S0025-5718-1978-0502013-4 [PDF]; doi:10.2307/2006350 [PDF], JSTOR; MR0502013 (80h:12004) (reviewed by S. P. Demuškin; reprinted in Reviews in Number Theory, volume 5A, R26-263) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0401.12008 (reviewed by H. C. Williams) [scanned image].

[40] C. D. Walter : Brauer’s class number relation, Acta Arith. XXXV (1979), fasc. 1, 33–40; aa3514.pdf, icm; MR0536878 (82d:12006) (reviewed by T. Lepistö; reprinted in Reviews in Number Theory, volume 5A, R26-274) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0339.12007 (Autorreferat) [scanned image].

[41] C. D. Walter : Kuroda’s class number relation, Acta Arith. XXXV (1979), fasc. 1, 41–51; aa3515.pdf, icm; MR0536879 (82d:12007) (reviewed by T. Lepistö; reprinted in Reviews in Number Theory, volume 5A, R26-274E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0339.12008 (Autorreferat) [scanned image].

[42] N. Moser : Unités et nombre de classes d’une extension galoisienne diédrale de Q, Abh. Math. Sem. Univ. Hamburg 48 (1979), 54–75; DOI: 10.1007/BF02941290 [9v0vu4272t785v31.pdf (925.4 KB)]; MR0537446 (81h:12009) (reviewed by H. Cohn; reprinted in Reviews in Number Theory, volume 5A, R22-229E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0387.12005 (Autorreferat) [scanned image].

[43] N. Moser : Sur les unités d’une extension galoisienne non abélienne de degré pq du corps des rationnels p et q nombres premiers impairs, Ann. Inst. Fourier (Grenoble) 29 (1979), fasc. 1, xv, 137–158;

                                                      doi: 10.5802/aif.729 [AIF_1979__29_1_137_0.pdf];
                                                      AIF_1979__29_1_137_0.djvu,
                                                      AIF_1979__29_1_137_0.pdf, Numdam; MR0526781

(80e:12010) (reviewed by P. Damey; reprinted in Reviews in Number Theory, volume 5A, R22-268) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0387.12003 (Autorreferat) [scanned image].

[44] C. J. Parry : A genus theory for quartic fields, J. Reine Angew. Math. 314 (1980), 40–71; doi: 10.1515/crll.1980.314.40 [PDF (2772 KB)];

PPN243919689_0314_log8.pdf (2,34 MB), GDZ;

[PDF], DigiZeitschriften; MR0555904 (81j:12002) (reviewed by G. Whaples; reprinted in Reviews in Number Theory, volume 5A, R16-226E) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0421.12006 (Autorreferat) [scanned image].

[45] N. Nakagoshi : On the class number relations of abelian extensions whose Galois groups are of type (p, p), Math. Rep. Toyama Univ. 4 (1981), no. 1, 91–106; MR0627966 (83b:12008) (reviewed by Y. Furuta; reprinted in Reviews in Number Theory, volume 5A, R20-245) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0482.12005 (reviewed by R. Gillard) [scanned image].

[46] 片岡俊孝 [Kataoka, T.] : Class number relations, P 進 L 関数と代数体の整数論 [P-Adic L-Functions and Algebraic Number Theory] (京都大学数理解析研究所, 1981) (森田康夫編), 数理解析研究所講究録 [RIMS Kokyuroku] 411, 1981, 142–161; 0411-12.pdf; 0411-12.pdf (994.14 kB), URI: http://hdl.handle.net/2433/102414.

[47] R. Bölling : Zur Klassenzahl nicht galoissches Körper in Diedererweiterungen über Q mit besonder Berucksichtungung kubischer Körper, I, Math. Nachr. 118 (1984), Issue 1, 271–284; DOI: 10.1002/mana.19841180120 [PDF (792 K)]; MR0773626 (87a:11111) (reviewed by N. Klingen; reprinted in Reviews in Number Theory, volume 5B, R29-676) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0562.12008 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[48] N. Nakagoshi : A note on l-class groups of certain algebraic number fields, J. Number Theory 19 (1984), no. 2, 140–147; doi:10.1016/0022-314X(84)90098-2 [PDF (337 K)]; MR0762764 (86b:11072) (reviewed by Y. Furuta; reprinted in Reviews in Number Theory, volume 5B, R29-656) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0548.12007 (reviewed by Ch. Parry) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[49] J. Cougnard : Groupe des unités et nombre de classes de certains extensions diédrales de degré 8 de Q, Théorie des Nombres, Besançon, 1983–1984, 21 pp., Publ. Math. Fac. Sci., Exp. No. 2, Besançon, Univ., Franche-Comté, Besançon, 1984; Cougnard.pdf; MR0803698 (87e:11126) (reviewed by C. J. Parry; reprinted in Reviews in Number Theory, volume 5B, R29-711) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0565.12005 (reviewed by H. Yokoi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[50] A. Endô : Class number relation between certain sextic number fields, Proc. Amer. Math. Soc. 95 (1985), no.

2, 199–204; DOI: 10.1090/S0002-9939-1985-0801323-6 [PDF]; doi:10.2307/2044512 [PDF], JSTOR; MR0801323 (86k:11062) (reviewed by R. Ernvall; reprinted in Reviews in Number Theory, volume 5B, R29-669) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0583.12003 (reviewed by C. Parry) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[51] T. Ono : On some class number relations for Galois extensions, Proc. Japan Acad. Ser. A Math. Sci. [日本学士院紀要] 61 (1985), No. 10, 311–312; doi:10.3792/pjaa.61.311, euclid.pja/1195514532 [PDF (176 KB)]; MR0834535 (87f:11089) (reviewed by J. S. Joel; reprinted in Reviews in Number Theory, volume 5B, R29-684) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0583.12005 (reviewed by K. Lei) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[52] T. Ono : On some class number relations for Galois extensions, Nagoya Math. J. 107 (1987), 121–133; euclid.nmj/1118780772 [PDF (923 KB)]; MR0909252 (89f:11153) (reviewed by R. T. Bumby; reprinted in Reviews in Number Theory, volume 5B, R29-743) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0642.12018 (reviewed by S. Kanemitsu) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[53] G. Cornell : Group theory and the class group, Number theory and applications. (Banff, AB, 1988) (R. A. Mollin, ed.), NATO Adv. Sci. Inst. Ser. C. Math. Phys. Sci., 265, Kluwer Acad. Publ., Dordrecht, 1989, pp. 347–352; MR1123080 (92g:11106) (reviewed by W. Narkiewicz; reprinted in Reviews in Number Theory, volume 5B, R29-834) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0698.12004 (reviewed by H. Opolka) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[54] K. Ohta : On the class numbers and the ideal class groups of certain algebraic number fields, Bull. Fac. Gen. Ed. Gifu Univ. [岐阜大学教養部研究報告] No. 25 (1989), 57–67 (1990); MR1049927 (91e:11127) (reviewed by H. Yokoi; reprinted in Reviews in Number Theory, volume 5B, R29-795) [html, PDF, DVI, PostScript, BibTeX].

[55] G. Cornell : A note on the class group of composita, J. Number Theory 39 (1991), no. 1, 1– 4; doi:10.1016/0022-314X(91)90028-A [PDF (230 K)]; MR1123163 (93b:11145a) (reviewed by J. W. Sands; reprinted in Reviews in Number Theory, volume 5B, R29-844a) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0735.11060 (reviewed by T. Takeuchi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[56] D. Z. Moroz : Letter to the editor: “A note on the class group of composita”, J. Number Theory 40 (1992), no. 1, 125; doi:10.1016/0022-314X(92)90033-L [PDF (26 K)]; MR1145859 (93b:11145b) (reviewed by J. W. Sands; reprinted in Reviews in Number Theory, volume 5B, R29-844b) [html, PDF, DVI, PostScript, BibTeX].

[57] R. I. Berger : Hasse’s class number product formula for generalized Dirichlet fields and other types of number fields, Manuscripta Math. 76 (1992), fasc. 3–4, 397–406; DOI: 10.1007/BF02567768 [y420155088532728.pdf

(512.8 KB)]; PPN365956996_0076_log31.pdf (969,71

KB), GDZ; [PDF], DigiZeitschriften; MR1185027 (93h:11121) (reviewed by S. Louboutin; reprinted in Reviews in Number Theory, volume 5B, R29-862) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0779.11051 (reviewed by M. Kolster) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[58] S. Louboutin : On the class number one problem for non-normal quartic CM-fields, Tôhoku Math. J.

(2) [東北數學雑誌第二輯] 46 (1994), no. 1, 1– 12; doi:10.2748/tmj/1178225798, euclid.tmj/1178225798 [PDF (947 KB)]; 46_1.pdf (461 K), Journal@rchive, JST; MR1256724 (94m:11130) (reviewed by K. Yamamura; reprinted in Reviews in Number Theory, volume 5B, R29-894) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0796.11050 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[59] F. Lemmermeyer : Kuroda’s class number formula, Acta Arith. LXVI (1994), fasc. 3, 245–260; aa6635.pdf, icm; MR1276992 (95f:11090) (reviewed by C. J. Parry; reprinted in Reviews in Number Theory, volume 5B, R29-908) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0807.11052 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[60] E. Kani : Discriminants of Hermitian R[G]-modules and Brauer’s class number relation, Algebra and number theory (Essen, 1992) (G. Frey and J. Ritter, eds.), de Gruyter, Berlin, 1994, pp. 43–135; [PDF (11,726.1 KB), PDF with Links (9,851.6 KB)]; MR1285363 (95h:11120) (reviewed by M. E. Keating; reprinted in Reviews in Number Theory, volume 5B, R29-911) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0811.11067 (reviewed by J. Ritter) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[61] F. Lemmermeyer : Ideal class groups of cyclotomic number fields I, Acta Arith. LXXII (1995), fasc. 4, 347–359; aa7245.pdf, icm; MR1348202 (96h:11111) (reviewed by by S. Louboutin; reprinted in Reviews in Number Theory, volume 5B, R18-692) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0837.11059 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[62] K. Yamamura : Maximal unramified extensions of imaginary quadratic number fields of small conductors, J. Théor. Nombres Bordeaux 9 (1997), Fasc. 2, 405–446;

                                                      doi: 10.5802/jtnb.211 [JTNB_1997__9_2_405_0.pdf];
                                                      JTNB_1997__9_2_405_0.djvu,  JTNB_1997__9_2_

405_0.pdf, Numdam; MR1617407 (99e:11139) (reviewed by F. Diaz y Diaz) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0905.11048 (reviewed by E. Benjamin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[63] F. Lemmermeyer : Ideal class groups of cyclotomic number fields II, Acta Arith. LXXXIV (1998), fasc. 1, 59– 70; aa8415.pdf, icm; MR1613302 (99a:11121) (reviewed by S. Louboutin) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0901.11031 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[64] W. Bosma and B. de Smit : Class number relations from a computational point of view, J. Symbolic Comput. 31 (2001), no. 1–2, 97–112; doi:10.1006/jsco.1999.1016 [PDF (353 K)]; MR1806209 (2002a:11144) (reviewed by A. B. Ekin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0982.11073 (reviewed by M. Pohst) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[65] F. Lemmermeyer : Galois action on class groups, J. Algebra 264 (2003), no. 2, 553–564; doi:10.1016/S0021-8693(03)00122-4 [PDF (127 K)]; MR1981421 (2004e: 11125) (reviewed by A. Mimouni) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1032.12008 (reviewed by C. Parry) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[66] F. Lemmermeyer : Class groups of dihedral extensions, Math. Nachr. 278 (2005), Issue 6, 679–691; DOI: 10.1002/mana.200310263 [PDF (204 K)]; MR2135500 (2006k:11217) (reviewed by M. R. Bush) [html, PDF, DVI, PostScript, BibTeX]; Zbl 1067.11069 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[67] 山村健 : (代数体の) Class number relation に関する文献 (2006 年 10 月更新版), 第 4 回北陸数論研究集会報告集 [Proceedings of the 4th Workshop on Number Theory] (金沢大学サテライトプラザ, 2005) (平林幹人, 野村明人, 木村巌, 菅野孝史編), 2006, pp. 182–188.